# Adaptive RKHS Fourier Features for Compositional Gaussian Process Models

### v1

@@ 3.2 Variational Fourier Response Features for LFMs
By the closure of GPs under linear operations, the output $f$ and the projection $\mathbf{v}$ share a joint augmented GP prior. The covariance matrix of inducing variables $\mathrm{Cov}[\mathbf{v},\mathbf{v}]$ has a low-rank-plus-diagonal structure if inducing frequencies $\mathbf{z}=[z_{m}]_{m=1}^{M}$ are harmonic on $[a,b]$, i.e., $z_{m}=\frac{2\pi m}{b-a}$, facilitating faster posterior computation (Hensman et al., 2018). For a given input $t$, the cross-covariance of the output process $f$ and the inducing variable $v_{m}$ is computed as
@@
$\mathrm{Cov}[f(t),v_{m}]=\int_{-\infty}^{t}G(t-\tau)\phi_{m}(\tau)\mathop{}\!\mathrm{d}\tau=\begin{cases}\frac{\cos(z_{i}(t-a)+\theta)}{\beta\sqrt{z_{i}^{2}+\gamma^{2}}}+\xi_{i}&i=0,1,\ldots,M,\\ \frac{\sin(z_{i}(t-a)+\theta)}{\beta\sqrt{z_{i}^{2}+\gamma^{2}}}+\xi_{i}&i=M+1,\ldots,2M,\end{cases}$ (12)
where $z_{0}=0$, cutoff frequency $\gamma=\frac{\alpha}{\beta}$, phase shift $\theta=-\arctan(\frac{z_{i}}{\gamma})$, and $\xi_{i}$ represents an exponential decay term. Since the integration variable $\tau$ ranges from negative infinity and the inner product $\langle k(\tau,\cdot),\phi(\cdot,z)\rangle_{\mathcal{H}}$ has different expressions beyond $\tau\in[a,b]$, the covariance of $f$ and $v_{m}$ for $t\in\mathbb{R}$ emerges as a continuous piece-wise function (see Fig. 1 right). The detailed derivation and complete expressions of $\text{Cov}[f(t),v(z)]$ for Matérn-$\tfrac{1}{2}/\frac{3}{2}/\frac{5}{2}$ are given in Appx. 10 and illustrated in Fig. 7 and 8.
The derived inter-domain features (12) reflect the filtering effect of the system, i.e., how the ODE adaptively transforms the frequency components of latent forces to the output through amplitude attenuation and phase shift. By analogy with the frequency response of linear systems, we name the obtained Fourier features from RKHS as Variational Fourier Response Features (VFRFs). Fig. 1, right, depicts the VFRFs of the LFM from the left subplot ($\gamma=4$, solid red) and the VFFs of the corresponding latent force (dash blue). They show that the VFRFs are learnable basis functions that adjust both the amplitude and the phase according to the input frequencies and the ODE parameters. Particularly, the system will allow nearly all frequency components of the input process to pass through as $\gamma\rightarrow+\infty$. Under this condition, VFRFs converge to the VFFs of inter-domain GPs. We would like to emphasise here that the features derived from (11) can apply to more general inter-domain GPs with other linear transformations $G(\cdot)$, not just limited to LFMs. Moreover, for stable dynamical systems governed by higher-order ODEs within the LFM framework, the derivation of VFRFs described above can be readily extended by designing properly scaled Green’s functions  [^2] The “stable” system here means its ODE has a general solution involving exponentially decaying terms, in other words, the roots of the characteristic equation lie in the left half of the complex plane. In this case, (11) may still produce sinusoidal features..
@@ 3.3 Deep LFMs with VFRFs
DLFMs extend the concept of shallow LFMs by stacking them in a non-parametric cascade, similar to DGPs. This hierarchical structure allows DLFMs to model the non-stationarities present in nonlinear dynamical systems. In this section, we detail the construction of a hierarchical composition of $L$ LFMs within the framework of variational DGPs, each governed by the modified ODE and enhanced with VFRFs for variational approximation of the posterior. We leverage the layer-wise Monte Carlo technique in doubly stochastic variational inference (Salimbeni and Deisenroth, 2017) to allow functional samples to propagate through the compositional architecture efficiently.
The first layer of a DLFM processes a $D^{0}$-dimensional input $\mathbf{t}=[t_{d}]_{d=1}^{D^{0}}$ and outputs a $D^{1}$-dimensional independent process $\boldsymbol{g}^{1}(\mathbf{t})=[g^{1}_{r}(\mathbf{t})]_{r=1}^{D^{1}}$ (the superscripts indicate the layer index). To extend the application of VFRFs to multidimensional inputs, we follow Hensman et al. (2018) to employ additive LFM kernels for each output dimension, i.e., each output dimension $g^{1}_{r}(\mathbf{t})$ is modelled as $g^{1}_{r}(\mathbf{t})=\sum_{d=1}^{D^{0}}f^{1}_{d}(t_{d})$, where $\{f^{1}_{d}\}_{d=1}^{D^{0}}$ are LFMs with ODE-induced covariance functions. In this work, we assume the LFMs $f^{1}_{d}$ are independent, but this assumption can be relaxed by allowing them to share the same latent forces, which can lead to more complex kernels for the outputs.
@@
$\mathrm{Cov}[g^{1}_{r}(\mathbf{t}),v^{1}_{m,d}]=\int_{-\infty}^{t_{d}}G^{1}_{d}(t_{d}-\tau)\phi_{m}(\tau)\mathop{}\!\mathrm{d}\tau,\quad\mathrm{Cov}[v^{1}_{m,d},v^{1}_{m^{\prime},d^{\prime}}]=0\ \ (d\neq d^{\prime}).$ (13)
Assuming a variational distribution $q(\mathbf{V}^{1})$, the approximate posterior $q(\boldsymbol{g}^{1}\mid\mathbf{t})$ of the first layer is derived by substituting $k(\mathbf{x},\mathbf{Z})$ and $\mathbf{K_{ZZ}}$ in (4) with the expressions of VFRFs. Samples from this approximate posterior are drawn using the re-parameterisation trick (Kingma et al., 2015).
@@
$\mathrm{ELBO}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\mathbb{E}_{q(\mathbf{g}^{L}_{i}\mid\mathbf{t}_{i})}\log\left[p(\mathbf{y}_{i}\mid\mathbf{g}^{L}_{i})\right]-\frac{1}{N}\sum_{l=1}^{L}\mathrm{KL}\left[q(\mathbf{V}^{l})\mid\mid p(\mathbf{V}^{l})\right],$ (14)
@@ 4 Related Work
LFMs presented a physically-inspired approach to combining data-driven modelling with differential equations (Alvarez et al., 2009). Álvarez et al. (2010) further proposed variational inducing functions to handle non-smooth latent processes within convolved GPs (Alvarez and Lawrence, 2011). Our model builds upon LFMs and DGPs (Salimbeni and Deisenroth, 2017). Recently, various approximate inference methods have been explored for DGP-based models, which are generally categorised into variational inference techniques (Salimbeni and Deisenroth, 2017; Salimbeni et al., 2019; Lindinger et al., 2020) and Monte Carlo approaches(Havasi et al., 2018).
As outlined in Section 1, RFFs (Rahimi and Recht, 2007) and VFFs (Hensman et al., 2018) have recently been incorporated into GP models. RFFs were used in shallow LFMs models to approximate covariance matrices (Guarnizo and Álvarez, 2018) and expanded to a deeper architecture (McDonald and Álvarez, 2021, 2023). While related to these studies, our approach primarily uses features similar to VFFs within the scope of inter-domain GPs (Van der Wilk et al., 2020). Unlike RFF-based DGP models, which often alter the original covariance functions by introducing a fully parametric variational distribution over random frequencies, our model maintains the integrity of the original kernel forms and approximates the DGP posterior directly. Another closely related work is inter-domain DGPs (IDDGPs) (Rudner et al., 2020), which employ fixed VFFs. In contrast, our model extends the compositional inter-domain GPs by integrating ODEs to provide trainable, physics-informed RKHS Fourier features. Appx. 7 illustrates a high-level comparison of our work with related works.
@@ 5 Experiments
This section presents experiments designed to evaluate our model using VFRFs. We begin by examining the approximation quality of shallow LFMs with VFRFs and RFFs on synthetic data. Additionally, we evaluate our model on a highly non-stationary speech signal dataset, as well as on benchmark regression tasks, comparing to a variety of baselines in both cases. We will make the implementation publicly available after the review period.
Posterior Approximation of Shallow LFMs
VFRFs and RFFs both leverage Fourier features to facilitate approximate inference in LFMs. In Fig. 2, we compare the approximation quality of VFRFs and RFFs in a regression task using models with a Matérn-$\frac{1}{2}$ kernel. The kernel’s parameters and the noise variance are initially optimised by maximising the marginal likelihood of an exact LFM, then fixed across all models. The frequencies of RFFs are sampled from the corresponding Cauchy distribution (detailed in Appx. 9.2). Fig. 2 shows that the model with 20 VFRFs fits the data points reasonably well. Increasing VFRFs to 80 fills in the details of the region with more observations, and the approximate predictive posterior is quite close to the exact one. The same number of RFFs yields a poorly fitted predictive mean and tends to underestimate the variance in different regions of the input space, a phenomenon known as variance starvation (Wang et al., 2018). Achieving a comparable approximation requires more RFFs due to the heavy-tailed spectral density of Matérn-$\frac{1}{2}$ kernels.
Figure 2: Illustrative example of Matérn-$\tfrac{1}{2}$ LFM posterior with VFRFs / RFFs. The model’s feature is indicated at the lower right. Top row: predictive posteriors of 20, 80, and 500 RFFs. Bottom row: predictive posteriors of 20 and 80 inducing frequencies and an exact LFM. Noisy observations are marked with red dots, posterior predictive means with blue lines, and uncertainty (one or two standard deviations) with varying shades of blue. In this example, VFRFs show a better approximation to the true posterior, whereas RFRFs indicate variance underestimation with fewer features.
@@
Synthetic Data for Deep Structures
To further evaluate our DLFM-VFRF’s performance against other models, we conduct tests on a synthetic multi-step function (Fig. 3). This task is challenging for shallow GP models due to the need to capture global structures in highly non-stationary data (Rudner et al., 2020). Although equipped with VFRFs, our single-layer LFM (upper left) struggles to fit the non-stationarity with a stationary kernel. In contrast, the models with compositional layers exhibit better performance. The DLFM-RFF (lower left) (McDonald and Álvarez, 2021) generates high-frequency, wiggly posterior predictive samples, resulting in an easily over-fitting model struggling to seize the slow-changing trend in the data. DGP-RBF (upper left), IDDGP (lower middle), and our DLFM-VFRF (lower right) offer smoother samples from the posterior distributions. Due to the VFRFs’ flexibility to capture the global data structure, our DLFM-VFRF outperforms both the DGP-RBF with local inducing points and the IDDGP, especially inside $[0.8,1]$. Our model provides a more accurate predictive mean throughout the steps and at abrupt step transitions, and demonstrates narrower confidence intervals, indicating a better uncertainty calibration. In the plot, GP models based on function-space inference tend to revert to prior distributions outside the observed data range, displaying wide uncertainty bands. In contrast, the DLFM-RFF yields relatively more confident non-zero predictions in these areas.
TIMIT Speech Signals
We apply our model to a regression task on the TIMIT dataset, a speech recognition resource previously used by Rudner et al. (2020), to explore the capability of GP-based models in handling complex, non-stationary data. The dataset features rapidly changing audio waves, posing significant challenges for shallow GP models reliant on local approximation. Initially, we apply a moving average filter to smooth the audio waves and select the first 10,000 data points for our analysis, reserving 30 % as test data. Our method uses the Matérn-$\frac{3}{2}$ kernels.
(a) Learning progression of DLFMs and IDDGPs
@@
Figure 4: (a) Learning progression of DLFMs and IDDGPs with $M$ inducing frequencies on the TIMIT dataset, presented in negative ELBO, test average RMSE and NMLL. The DLFM in yellow maintains fixed $\beta=10^{-6}$ throughout the first 2000 training epochs, after which $\alpha/\beta$ are allowed to vary. The DLFM in red employs trainable ODE parameters from the start. The DLFMs with VFRFs consistently outperform the IDDGPs; (b) Mean standardised RMSE and NMLL with the standard deviations (over 10 random seeds) for models employing varying numbers of inducing frequencies. The numbers following the hyphen in the y-axis labels indicate the number of inducing frequencies/points. A lower value (to the left) indicates better performance.

### v2

@@ 3.2 Variational Fourier Response Features
By the closure of GPs under linear operations, the output $f$ and the projection $\mathbf{v}$ share a joint augmented GP prior. The covariance matrix of inducing variables $\mathrm{Cov}[\mathbf{v},\mathbf{v}]$ has a low-rank-plus-diagonal structure if inducing frequencies $\mathbf{z}=[z_{m}]_{m=1}^{M}$ are harmonic on $[a,b]$, i.e., $z_{m}=\frac{2\pi m}{b-a}$, facilitating faster posterior computation (Hensman et al.,, 2018). For a given input $t$, the cross-covariance of the output process $f$ and the inducing variable $v_{m}$ is computed as
@@
$\mathrm{Co}$ $\mathrm{v}[f(t),v_{m}]=\int_{-\infty}^{t}G(t-\tau)\phi_{m}(\tau)\mathop{}\!\mathrm{d}\tau$ (10)
=$=\begin{cases}\frac{\cos(z_{i}(t-a)+\theta)}{\beta\sqrt{z_{i}^{2}+\gamma^{2}}}+\xi_{i}&i=0,\ldots,M,\\ \frac{\sin(z_{i}(t-a)+\theta)}{\beta\sqrt{z_{i}^{2}+\gamma^{2}}}+\xi_{i}&i=M+1,\ldots,2M,\end{cases}$
where $z_{0}=0$, cutoff frequency $\gamma=\frac{\alpha}{\beta}$, phase shift $\theta=-\arctan(\frac{z_{i}}{\gamma})$, and $\xi_{i}$ represents an exponential decay term. Since the integration variable $\tau$ ranges from negative infinity and the inner product $\langle k(\tau,\cdot),\phi(\cdot,z)\rangle_{\mathcal{H}}$ has different expressions beyond $\tau\in[a,b]$, the covariance of $f$ and $v_{m}$ for $t\in\mathbb{R}$ emerges as a continuous piece-wise function (see Fig. 1 right). The detailed derivation and complete expressions of $\text{Cov}[f(t),v(z)]$ for Matérn-$\tfrac{1}{2}/\frac{3}{2}/\frac{5}{2}$ are given in Appendix D and illustrated in Fig. 8 and 9.
The derived inter-domain features (10) reflect the filtering effect of the system, i.e., how the ODE adaptively transforms the frequency components of latent forces to the output through amplitude attenuation and phase shift. By analogy with the frequency response of linear systems, we name the obtained Fourier features from RKHS as Variational Fourier Response Features (VFRFs). Fig. 1, right, depicts the VFRFs of the LFM from the left subplot ($\gamma=4$, solid red) and the VFFs of the corresponding latent force (dash blue). They show that the VFRFs are learnable basis functions that adjust both the amplitude and the phase according to the input frequencies and the ODE parameters. Particularly, the system will allow nearly all frequency components of the input process to pass through as $\gamma\rightarrow+\infty$. Under this condition, the VFRFs converge to the VFFs. We would like to emphasise here that the features derived from (9) can apply to more general inter-domain GPs with other linear transformations $G(\cdot)$, not just limited to LFMs. Moreover, for stable dynamical systems governed by higher-order ODEs within the LFM framework, the derivation of the VFRFs described above can be readily extended by using corresponding Green’s functions.
@@ 3.3 Deep LFMs with VFRFs
DLFMs extend the concept of shallow LFMs by stacking them in a non-parametric cascade, similar to DGPs. This hierarchical structure allows DLFMs to model the non-stationarities present in nonlinear dynamical systems. In this section, we detail the construction of a hierarchical composition of $L$ LFMs within the framework of variational DGPs, each governed by the modified ODE and enhanced with VFRFs for variational approximation of the posterior. Fig. 2 gives a conceptual illustration of how our proposed DLFM differs from a DGP. We leverage the layer-wise Monte Carlo technique in doubly stochastic variational inference (Salimbeni and Deisenroth,, 2017) to allow functional samples to propagate through the compositional architecture efficiently.
The first layer of a DLFM processes a $D^{0}$-dimensional input $\mathbf{t}=[t_{d}]_{d=1}^{D^{0}}$ and outputs a $D^{1}$-dimensional independent process $\boldsymbol{g}^{1}(\mathbf{t})=[g^{1}_{r}(\mathbf{t})]_{r=1}^{D^{1}}$ (the superscripts indicate the layer index). To extend the application of VFRFs to multidimensional inputs, we follow Hensman et al., (2018) to employ additive LFM kernels for each output dimension, i.e., each output dimension $g^{1}_{r}(\mathbf{t})$ is modelled as $g^{1}_{r}(\mathbf{t})=\sum_{d=1}^{D^{0}}f^{1}_{d}(t_{d})$, where $\{f^{1}_{d}\}_{d=1}^{D^{0}}$ are LFMs with ODE-induced covariance functions. In this work, we assume the LFMs $f^{1}_{d}$ are independent, but this assumption can be relaxed by allowing them to share the same latent forces, which can lead to more complex kernels for the outputs.
@@
$\mathrm{Cov}[g^{1}_{r}(\mathbf{t}),v^{1}_{m,d}]$ $=\int_{-\infty}^{t_{d}}G^{1}_{d}(t_{d}-\tau)\phi_{m}(\tau)\mathop{}\!\mathrm{d}\tau,$
$\mathrm{Cov}[v^{1}_{m,d},v^{1}_{m^{\prime},d^{\prime}}]$ $=0\ \ (d\neq d^{\prime}).$
Assuming a variational distribution $q(\mathbf{V}^{1})$, the approximate posterior $q(\boldsymbol{g}^{1}\mid\mathbf{t})$ of the first layer is derived by substituting $k(\mathbf{x},\mathbf{Z})$ and $\mathbf{K_{ZZ}}$ in (1) with the expressions of VFRFs. Samples from this approximate posterior are drawn using the re-parameterisation trick (Kingma et al.,, 2015).
@@
$\mathrm{ELBO}=$ $\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\mathbb{E}_{q(\mathbf{g}^{L}_{i}\mid\mathbf{t}_{i})}\log\left[p(\mathbf{y}_{i}\mid\mathbf{g}^{L}_{i})\right]$
$-\frac{1}{N}\sum_{l=1}^{L}\mathrm{KL}\left[q(\mathbf{V}^{l})\mid\mid p(\mathbf{V}^{l})\right],$
@@ 4 RELATED WORK
LFMs present a physically-inspired approach to combining data-driven modelling with differential equations (Alvarez et al.,, 2009). Álvarez et al., (2010) further proposed variational inducing functions to handle non-smooth latent processes within convolved GPs (Alvarez and Lawrence,, 2011). Our model builds upon LFMs and DGPs (Salimbeni and Deisenroth,, 2017). Recently, various approximate inference methods have been explored for DGP-based models, which are generally categorised into variational inference techniques (Salimbeni and Deisenroth,, 2017; Salimbeni et al.,, 2019; Lindinger et al.,, 2020) and Monte Carlo approaches(Havasi et al.,, 2018).
As outlined in Section 1, RFFs (Rahimi and Recht,, 2007) and VFFs (Hensman et al.,, 2018) have recently been incorporated into GP models. RFFs were used in shallow LFMs models to approximate covariance matrices (Guarnizo and Álvarez,, 2018) and expanded to a deeper architecture (McDonald and Álvarez,, 2021, 2023). VFFs were once integrated with harmonizable mixture kernels in shallow GP models (Shen et al.,, 2019). While related to these studies, our approach primarily uses features similar to VFFs within the scope of inter-domain GPs (Lázaro-Gredilla and Figueiras-Vidal, 2009a, ; Van der Wilk et al.,, 2020). Unlike RFF-based DGP models, which often modify the original covariance functions by introducing a fully parametric variational distribution over random frequencies, our model preserves the integrity of the original kernel forms and approximates the DGP posterior directly. Another closely related work is Inter-Domain DGPs (IDDGPs) (Rudner et al.,, 2020), which employ fixed VFFs without ODEs. We provide an illustrative plot of IDDGP in Fig. 2(a). In contrast, our model extends the compositional inter-domain GPs by integrating ODEs to provide trainable, physics-informed RKHS Fourier features. Fig. 7 in Appendix A illustrates a high-level comparison of our work with related studies.
@@ 5 EXPERIMENTS
This section presents experiments designed to evaluate our model using VFRFs. We begin by examining the approximation quality of shallow LFMs with VFRFs and RFFs on synthetic data. We then evaluate our model on a highly non-stationary speech signal dataset and benchmark regression tasks, comparing it to various baselines in both cases. [^1] Our code is publicly available in the repository: https://github.com/shixinxing/LFM-VFF
Figure 3: Illustrative example of Matérn-$\tfrac{1}{2}$ LFM posteriors with VFRFs / RFFs. The model’s feature is indicated at the lower right. Top row: predictive posteriors of 20, 80, and 500 RFFs. Bottom row: predictive posteriors of 20 and 80 inducing frequencies and an exact LFM. Noisy observations are marked with red dots, posterior predictive means with blue lines, and uncertainty (one or two standard deviations) with varying shades of blue. In this example, VFRFs show a better approximation to the true posterior, whereas RFFs indicate variance underestimation with fewer features.
@@
5.1 Synthetic Datasets
We first evaluate the shallow LFMs and DLFMs using the proposed VFRFs on two synthetic datasets, respectively.
5.1.1 Posterior Approximation for Shallow LFMs
VFRFs and RFFs both leverage Fourier features to facilitate approximate inference in LFMs. In Fig. 3, we compare the approximation quality of VFRFs and RFFs in a regression task using models with a Matérn-$\frac{1}{2}$ kernel. The kernel’s parameters and the noise variance are initially optimised by maximising the marginal likelihood of an exact LFM and then fixed across all models. The frequencies of RFFs are sampled from the corresponding Cauchy distribution of the kernel (detailed in Appendix C.2).
Fig. 3 shows that the model using 20 VFRFs has already fitted the data points reasonably well. Increasing VFRFs to 80 fills in the details of the region with more observations, and the approximate predictive posterior is quite close to the exact one. The same number of RFFs yields a poorly fitted predictive mean and tends to underestimate the variance in different regions of the input space, which is a phenomenon known as variance starvation (Wang et al.,, 2018). From the top row of Fig. 3, we can observe that achieving a comparable approximation requires more RFFs than VFRFs due to the heavy-tailed spectral density of the Matérn-$\frac{1}{2}$ kernels used in RFFs.
(a) Learning progression of DLFM-VFRF and IDDGP
@@ 5.1.1 Posterior Approximation for Shallow LFMs
Figure 5: (a) Learning progression of DLFMs and IDDGPs with $M$ inducing frequencies on the TIMIT dataset, presented in negative ELBO, test average RMSE and NMLL. The DLFM in yellow maintains fixed $\beta=10^{-6}$ throughout the first 2000 training iterations, after which $\alpha/\beta$ are allowed to vary. The DLFMs in red employ trainable ODE parameters from the start. The DLFM-VFRFs consistently outperform the IDDGPs; (b) Mean standardised RMSE and NMLL with the standard deviations (over 10 random seeds) for models employing varying numbers of inducing frequencies. The numbers following the hyphen in the y-axis labels indicate the number of inducing frequencies/points. A lower value (to the left) indicates better performance.
5.1.2 Multi-step Function for Deep Structures
To further evaluate our DLFM-VFRF’s performance against other models, we conduct tests on a synthetic multi-step function (as shown in Fig. 4). This task is challenging for shallow GP models due to the need to capture global structures in highly non-stationary data (Rudner et al.,, 2020).
Although equipped with VFRFs, our single-layer LFM (upper left) struggles to fit the non-stationarity with a stationary kernel. In contrast, the models with compositional layers exhibit better performance. The DLFM-RFF (lower left) (McDonald and Álvarez,, 2021) generates high-frequency, wiggly posterior predictive samples, resulting in an easily over-fitting model struggling to seize the slow-changing trend in the data. DGP-RBF (upper left), IDDGP (lower middle), and our DLFM-VFRF (lower right) offer smoother samples from the posterior distributions. Due to the VFRFs’ flexibility to capture the global data structure, our DLFM-VFRF outperforms both the DGP-RBF with local inducing points and the IDDGP, especially inside $[0.8,1]$. Our model provides a more accurate predictive mean throughout the steps and at abrupt step transitions and demonstrates narrower confidence intervals, indicating a better uncertainty calibration. In the plot, GP models based on function-space inference tend to revert to prior distributions outside the observed data range, displaying wide uncertainty bands. In contrast, the DLFM-RFF yields relatively more confident non-zero predictions in these areas.
To measure the performance quantitatively, we conducted additional experiments to train five independent copies of the IDDGP and our model. We summarise the Root Mean Square Error (RMSE) and Negative Marginal Log-Likelihood (NMLL) on the test points in the following Table 1. Additionally, the outputs of the intermediate layers for DGP, IDDGP, and our model are shown in Fig. 10 in Appendix E.
Table 1: Performance of IDDGP and our DLFM-VFRF on fitting the multi-step function over five runs (lower is better).
Model RMSE NMLL
IDDGP 0.107 $\pm$ 0.015 -1.051 $\pm$ 0.131
Ours 0.095 $\pm$ 0.010 -1.304 $\pm$ 0.120
5.2 TIMIT Speech Signals
We apply our model to a regression task on the TIMIT dataset, a speech recognition resource previously used by Rudner et al., (2020), to explore the capability of GP-based models in handling complex, non-stationary data. The dataset features rapidly changing audio waves, posing significant challenges for shallow GP models reliant on local approximation. Initially, we apply a moving average filter to smooth the audio waves and select the first 10,000 data points for our analysis, reserving 30% as test data. Our method uses the Matérn-$\frac{3}{2}$ kernels.